OpenStudy (lilai3):

x to the 7th pwer times x to the 5th power

OpenStudy (lgbasallote):

\[x^7 \times x^5\] use the rule \[a^m \times a^n = a^{mn}\]

OpenStudy (lilai3):

is it x to the 12th pwer

OpenStudy (lgbasallote):

yup

OpenStudy (lgbasallote):

i wrote that rule wrong lol

OpenStudy (lgbasallote):

\[\large a^m \times a^n = a^{m+ n}\]

OpenStudy (lilai3):

what is x to the 7th power times x to the megatice 3 power..and it's okay i'm just checking my hmwk

OpenStudy (lgbasallote):

what do you think?

OpenStudy (lilai3):

one over x to the seventh power times x to the third pwer

OpenStudy (lgbasallote):

is that your answer?

OpenStudy (lilai3):

yeah